A charge emanates an electric field of 7 N/C. If this field is placed on a rhombus surface that has a side 

length of 5mm and one of its angles is 60o

and is angled 60o from the surface, what is the electric flux?


1
Expert's answer
2021-02-11T08:54:59-0500

Find area of the rhombus:


"A=d^2 \text{ sin}\theta."


The flux is


"\Phi=EA\text{ sin}\theta=E(d\text{ sin}\theta)^2=1.31\cdot10^{-4}\text{ N m}^2/\text{ C}."